# ControlDreamer: Stylized 3D Generation with Multi-View ControlNet

### v2

@@ 4.4.2 MV-ControlNet as a plug-in for prior methods
Additionally, we demonstrate a scenario where MV-ControlNet is used in the refinement stage of existing methodologies, with the identical geometry and style prompts. DreamFusion [29] is employed to create the geometry model. We then compare the performance of MV-ControlNet against Magic3D [21]. As shown in Fig. 8 (b), our approach enhances 3D refinement and effectively avoids semantic drift, highlighting MV-ControlNet’s adaptability as a plug-in component across various methods.
5 Conclusions
We introduce ControlDreamer, a novel pipeline integrating the MV-ControlNet, our novel depth-aware multi-view diffusion model, to enable efficient, text-guided creation of stylized 3D models. This integration effectively generate stylized 3D models by aligning diverse geometries and styles given by textual descriptions. Additionally, our extensive benchmark for 3D style editing, covering a broad range of subjects, establishes a new standard in the field. Comparative analysis confirms that ControlDreamer outperforms existing text-to-3D methods in stylizing 3D models.
6 Limitations and Future Works
While ControlDreamer excels in generating stylized 3D models, it faces several limitations. Errors in pre-trained depth estimators can cause hole artifacts from the geometry stage to persist in the style stage. Another constraint is our experiment’s 256x256 resolution limit, due to MVDream’s training at this resolution. Integrating models like SDXL [28], trained on 1024x1024 images, could be a solution for higher-resolution outcomes. Future work will likely focus on developing an image-conditioned MV-ControlNet and exploring data scaling strategies.
Appendix 0.A Additional Experiments
0.A.1 Qualitative results
@@
In Fig. 10, we visualize more diverse qualitative results used for evaluating directional CLIP similarities and human evaluations.
0.A.2 Ablation Studies
0.A.2.1 Exploring 3D Representations in Style Stages
@@
Fig. 11 illustrates the approach of refining 3D models with NeRF instead of DMTet during the style stage. We observe that volume rendering with NeRF often leads to substantial alterations in the overall geometry, especially when geometry and style texts are semantically different, resulting in instability in depth-aware score distillation. Conversely, employing DMTet, coupled with our timestep annealing approach, proves more effective. It ensures appropriate alterations in the overall texture while maintaining the geometry aligned with the source mesh, leading to more seamless style generation. For NeRF-based 3D stylization, we sample diffusion timesteps from a range of [0.02, 0.5] to minimize significant geometry changes. MV-ControlNet often ignores camera conditions, leading to the creation of multiple artifacts, which further supports the use of DMTet in the style stage.
0.A.2.2 Advantages of MV-ControlNet in multi-view image generation
@@
Fig. 12 illustrates the advantages of the proposed MV-ControlNet through 2D image generation. In Fig. 12 (A), conventional methods exhibit significant semantic variations depending on random seed and camera variations, highly deviating from the original identity of the source image. In contrast, our proposed method demonstrates a stronger alignment with the depth condition, ensuring consistent object-centric generation quality even with different camera parameters and random seeds.
@@
In Fig. 12 (B), when providing text with complex concepts, as shown in (b), the baseline method, likewise as we explained in Fig. 1, displays stereotypical and geometry-biased generation which hampers the image generation quality. On the other hand, our method consistently generates high-quality images in line with the given text without forgetting important identities from the source (e.g., the backpack of an astronaut).
@@
In summary, the proposed MV-ControlNet is more strongly grounded to depth conditions and adept at reflecting complex concepts of the text for the generation of multi-view images.
0.A.3 Effectiveness of Prompt Engineering for 3D Generation
@@
When we use user-defined text prompts not included in the original training set of the MVDream model, we observe artifact generation in the learned source geometry, as shown in Fig. 13. However, we find that careful prompt engineering helps mitigate these artifacts. An example of such a refined text in our experiment is ‘Hulk, muscular, green, 4K, photorealistic’.
0.A.4 Geometry bias in Objaverse dataset
@@
We include additional results on the geometry bias discussed in the introduction of the main paper. Fig. 14 showcases text-based 3D model retrieval on Objaverse dataset, providing a straightforward visualization of the frequency of certain geometries. Notably, we show in the main text that 3D models predominantly associated with a shield are frequently generated from the prompt ‘Captain America’.
0.A.5 Enhanced Stylization in the Image-to-3D Framework
@@
Our methodology exhibits exceptional stylization capabilities, independent of the initialized geometry. Building on this strength, we extend our experiments to an image-to-3D framework [41] to evaluate its editing capabilities. On 3D models generated by ImageDream  [41], we modify the style with our MV-ControlNet. The results, as depicted in Fig. 15, highlight the adaptability of our approach to the image-to-3D conversion process.
0.A.6 Multi-view image editing
@@
In Fig. 16, we demonstrate the editing results of multi-view images using various prompts. For instance, with ’Superman’ as the geometry prompt, we add a cloak to the initial geometry. While the Prompt-to-Prompt (P2P) approach [12] relies on cross-attention control, our method incorporates depth conditions as an additional input to better align with the target style. P2P struggles to produce photorealistic images for counterfactual concepts like Spiderman wearing a cloak. In contrast, our method consistently achieves high-quality results with faithful text-alignments.
0.A.7 Failure cases
@@
In Fig. 17, we present examples of our failure cases. We empirically observe that when the style prompt significantly differs from the initial geometry, our method can struggle to generate visually appealing 3D models. This limitation likely stems from the insufficient capability of the pre-trained diffusion model, namely MVDream, to compose two distinct concepts.
Appendix 0.B Experiment Details
0.B.1 Baselines
@@
We utilize threestudio [10] implementation of Magic3D [21], Fantasia3D [5], and ProlificDreamer [42] as our baselines. Similar to ControlDreamer, all baselines undergo a refinement process on textured meshes derived from the pre-trained MVDream [39] using the DMTet [38] algorithm. Fantasia3D comprises two stages: geometry modeling and appearance modeling, and we use appearance modeling to modify the textured mesh. In the case of the ProlificDreamer, we employ a textured mesh fine-tuning stage for stylization. Additionally, despite ProlificDreamer not utilizing a separate shading model, we apply shading to ensure fair comparisons with other baselines. Our paired geometry and style prompts are shown in Table 3.
@@
0.B.2 Generated images used for MV-ControlNet training
In Fig. 18 and Fig. 19, we present additional examples used in training our MV-ControlNet. Both figures illustrate samples created through random sampling, with the camera views for each sample also randomized. These figures clearly show that samples with filtered $100K$ text corpus include a wide range of 3D assets, such as objects, animals, and characters. We utilized these high-quality multi-view images to train MV-ControlNet and thoroughly assessed its performance using diverse user-crafted prompts.
0.B.3 Human evaluations
Our human evaluation spans paired prompts from five domains, each involving six prompts for five methodologies: Fantasia3D, Magic3D, ProlificDreamer, MVDream, and our approach. Evaluations focus on editability and text alignment, with specific questions outlined in the table below.
Your task is to evaluate the quality of 4.0-second-long food video clips. Each clip showcases the results of editing 3D models using text prompts. Please choose the most favorable video based on the following criteria: 1) Overall editing quantity: Has the edited video completely changed in visual appearance compared to the Source? 2) Textual alignment: How well does the text description align with the edited video?
Figure 10: Additional qualitative results. We generate multiple styles on top of various source 3D models.
Figure 12: (A) The first column shows source multi-view images along with their depths. We showcase a comparison between images generated by MVDream (second column) and MV-ControlNet (third column), each originating from two different initial noises. MV-ControlNet’s depth-aware generation consistently preserves the source’s geometry, effectively handling variations in initial noise. (B) Our model excels in generating unique images, such as a knight with a backpack, and manages counterfactual prompts like ‘climber zebra’. It produces high-quality images, outperforming MVDream, which struggles with such tasks.
Figure 13: The comparison shows 3D models generated from user-provided texts versus ChatGPT-recommended ones. The left column has models from user texts; the right, those recommended by ChatGPT. Notably, user texts may cause artifacts, as seen in normal maps, whereas ChatGPT’s refined texts result in higher-quality 3D models.
Figure 15: Visualization of our stylization process within the image-to-3D pipeline. The source image prompts chosen for this demonstration were ‘pig’ and ‘crab’. Across these distinct subjects, our method maintained robust editing performance, showcasing its consistency in 3D stylization.
Figure 16: The comparative results on multi-view image editings. Our method consistently generates more photorealistic results than P2P [12].

### v3

@@ 4.4.2 MV-ControlNet as a plug-in for prior methods
Additionally, we demonstrate a scenario where MV-ControlNet is used in the refinement stage of existing methodologies, with identical geometry and style prompts. DreamFusion [Poole et al.(2023)Poole, Jain, Barron, and Mildenhall] is employed to create the geometry model. We then compare the performance of MV-ControlNet against Magic3D [Lin et al.(2023)Lin, Gao, Tang, Takikawa, Zeng, Huang, Kreis, Fidler, Liu, and Lin]. Fig. 5 (d) highlights MV-ControlNet’s adaptability as a plug-in component across various methods.
5 Discussion
We introduce ControlDreamer, a two-stage pipeline to improve text alignment in text-to-3D by decomposing the text into geometry and style prompts. To blend the geometry and style, we train MV-ControlNet, a depth-aware multi-view diffusion model, to enhance the geometry understanding during score distillation in the second stage. This integration effectively generates stylized 3D models by aligning diverse geometries and styles given by textual descriptions. Comparative results and theoretical analysis confirm that ControlDreamer outperforms existing text-to-3D methods in stylizing 3D models. Additionally, through comparisons with normal map and canny edge variants of MV-ControlNet, we demonstrated that depth conditioning is the most suitable for blending geometry and style.
While we improved text alignment in text-to-3D generation, some limitations remain. First, our training on single-object datasets limited our ability to experiment with complex scenes typical in 3D reconstruction. If multi-view diffusion models like ReconFusion [Wu et al.(2024)Wu, Mildenhall, Henzler, Park, Gao, Watson, Srinivasan, Verbin, Barron, Poole, et al.], trained on complex scenes, become publicly available, we could better assess ControlDreamer’s effectiveness. Additionally, though not our primary focus, future research could explore on various NeRF resolutions as these multi-view diffusion models advance to multiple resolutions and aspect ratios, similar to recent 2D image generative models [Podell et al.(2023)Podell, English, Lacey, Blattmann, Dockhorn, Müller, Penna, and Rombach].
Acknowledgement
This research was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) through a grant from the Korea government (MSIT) [No. RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)], the National Research Foundation of Korea (NRF) grant from the Korea government (MSIT) (No. 2022R1A3B1077720 and No. 2022R1A5A7083908), and the BK21 FOUR program for the Education and Research Program for Future ICT Pioneers at Seoul National University in 2024.
Appendix A Additional Experiments
A.1 Qualitative results
@@
In Figs. 7 and 8, we present a variety of qualitative results used for evaluating directional CLIP similarities and conducting human evaluations.
A.2 Ablation Studies
A.2.1 Exploring 3D Representations in Style Stages
@@
Fig. 9 demonstrates the approach of refining 3D models using NeRF instead of DMTet during the style stage. Volume rendering with NeRF often results in significant alterations to the overall geometry, particularly when geometry and style texts are semantically different, leading to instability in depth-aware score distillation. Conversely, using DMTet, coupled with our timestep annealing approach, is more effective. It ensures appropriate alterations in the overall texture while maintaining the geometry aligned with the source mesh, leading to more seamless style generation. For NeRF-based 3D stylization, we sample diffusion timesteps from a range of [0.02, 0.5] to minimize significant geometry changes. MV-ControlNet’s tendency to overlook camera conditions often leads to multiple artifacts, further substantiating the advantage of using DMTet in the style stage.
Figure 7: Our two-stage 3D generation pipeline first involves generating a coarse-grained geometry, followed by the generation of a fine-grained stylized 3D model using a style prompt.
Figure 8: Additional qualitative results. We generate a variety of styles on a range of source 3D models, demonstrating the versatility of our method.
Figure 10: (A) The first column displays source multi-view images along with their corresponding depths. We present a comparison of images generated by MVDream in the second column and MV-ControlNet in the third column, originating from two distinct initial noises. MV-ControlNet’s depth-aware generation consistently preserves the source’s geometry and effectively manages variations in initial noise. (B) Our model excels at generating unique images, such as a knight equipped with a backpack, and adeptly handles imaginative prompts like ‘climber zebra’. It consistently produces high-quality images, outperforming MVDream, which often struggles with such complex tasks.
A.2.2 Advantages of MV-ControlNet in multi-view image generation
@@
Fig. 10 illustrates the advantages of the proposed MV-ControlNet in 2D image generation. In panel (A), conventional methods exhibit significant semantic variations due to random seed and camera parameter changes, often deviating from the original identity of the source image. In contrast, our proposed method demonstrates stronger alignment with the depth condition, ensuring consistent, object-centric generation quality across different camera parameters and random seeds.
@@
In panel (B), when handling text with complex concepts, the baseline method displays stereotypical and geometry-biased generations that compromise image quality, as discussed earlier in Fig. 1 of the main text. Conversely, our method consistently produces high-quality images that align closely with the provided text, preserving key identities from the source, such as the backpack of an astronaut.
@@
In summary, MV-ControlNet is firmly grounded in depth conditions and adept at incorporating complex textual concepts into the generation of multi-view images.
A.3 Effectiveness of Prompt Engineering for 3D Generation
@@
When using user-defined text prompts that were not included in the training set of the MVDream model, we observe the generation of artifacts in the learned source geometry, as illustrated in Fig. 12. However, we find that careful prompt engineering can effectively mitigate these artifacts. An example of such refined text in our experiments is ‘Hulk, muscular, green, 4K, photorealistic’.
A.4 Geometry bias in Objaverse dataset
@@
We include further results on the geometry bias. Fig. 11 provides a straightforward visualization of text-based 3D model retrieval on the Objaverse dataset, showcasing the frequency of certain geometries. Notably, as we highlight in the main text, 3D models predominantly associated with a shield are frequently generated from the prompt ‘Captain America’.
A.5 Enhanced Stylization in the Image-to-3D Framework
@@
Our methodology demonstrates exceptional stylization capabilities, independent of the initialized geometry. Building on this strength, we have extended our experiments to include the image-to-3D framework as described in ImageDream [Wang and Shi(2023)], to further evaluate its editing capabilities. On 3D models generated using ImageDream, we applied our MV-ControlNet to modify the style. The results, depicted in Fig. 14, show the adaptability of our approach to the image-to-3D generation.
Figure 12: The comparison showcases 3D models generated from user-provided texts versus those recommended by ChatGPT. Notably, models from user texts may exhibit artifacts, particularly in normal maps, whereas models from ChatGPT’s refined texts generally result in higher-quality 3D models.
Figure 13: The comparative results on multi-view image editings. Our method consistently generates more photorealistic results than P2P [Hertz et al.(2023)Hertz, Mokady, Tenenbaum, Aberman, Pritch, and Cohen-Or].
A.6 Multi-view image editing
Prompt-to-Prompt (P2P) [Hertz et al.(2023)Hertz, Mokady, Tenenbaum, Aberman, Pritch, and Cohen-Or] is an image editing method that aligns the geometries of source and target images by injecting attention maps into diffusion models. For our qualitative comparison in the main text, injection ratios were set at 0.4 for cross-attention and 0.8 for self-attention to ensure strong geometric alignment. However, P2P is somewhat constrained by the requirement for equal token lengths in both source and target texts, which often results in less aligned text-guided editing outcomes. Conversely, our proposed MV-ControlNet, which utilizes depth conditions to maintain source geometry, is not restricted by text length, allowing for more creative and diverse prompts in the style stage.
@@ A.6 Multi-view image editing
In Fig. 13, we showcase editing results of multi-view images using various prompts. Using ‘Superman’ as a geometry prompt, we added a cloak to the initial geometry. The P2P method, which relies on cross-attention control, struggles with photorealistic images for complex scenarios like Spiderman wearing a cloak. Conversely, our MV-ControlNet incorporates depth conditions, achieving better style alignment and consistently producing high-quality, text-aligned images.
A.7 Failure cases
@@
In Fig. 15, we present examples of our failure cases. We empirically observe that when the style prompt significantly deviates from the initial geometry, our method struggles to generate visually appealing 3D models. This limitation likely arises from the pre-trained MVDream, which has insufficient capabilities to seamlessly integrate two distinct concepts.
Appendix B Experiment Details
B.1 Baselines
@@
We utilize the threestudio [Guo et al.(2023)Guo, Liu, Shao, Laforte, Voleti, Luo, Chen, Zou, Wang, Cao, and Zhang] implementation of Magic3D [Lin et al.(2023)Lin, Gao, Tang, Takikawa, Zeng, Huang, Kreis, Fidler, Liu, and Lin], Fantasia3D [Chen et al.(2023b)Chen, Chen, Jiao, and Jia], and ProlificDreamer [Wang et al.(2023)Wang, Lu, Wang, Bao, Li, Su, and Zhu] as our baselines. Similar to ControlDreamer, all baselines involve a refinement process on textured meshes derived from the pre-trained MVDream [Shi et al.(2023)Shi, Wang, Ye, Long, Li, and Yang] using the DMTet [Shen et al.(2021)Shen, Gao, Yin, Liu, and Fidler] algorithm. Specifically, Fantasia3D employs a two-stage process, focusing on appearance modeling in the second stage to modify the textured mesh. For ProlificDreamer, we engage in a textured mesh fine-tuning stage for stylization. Additionally, although ProlificDreamer does not utilize a separate shading model, we apply shading to ensure fair comparisons with other baselines. Our paired geometry and style prompts are detailed in Table 3.
@@
B.2 Generated images used for MV-ControlNet training
In Fig. 16 and Fig. 17, we present additional examples used in training our MV-ControlNet. Both figures illustrate samples created through random sampling, with the camera views for each sample also randomized. These figures clearly show that samples from a filtered $100K$ text corpus feature a wide range of 3D assets, such as objects, animals, and characters. We used these high-quality multi-view images to train MV-ControlNet.
B.3 Human evaluations
Our human evaluation spans paired prompts from five domains, with each domain involving six prompts across five methodologies: Fantasia3D, Magic3D, ProlificDreamer, MVDream, and our approach. The evaluations assess editability and text alignment, as detailed in the table below.
Human Evaluation Template Your task is to evaluate the quality of 4.0-second-long food video clips. Each clip showcases the results of editing 3D models using text prompts. Please choose the most favorable video based on the following criteria: 1) Overall editing quantity: Has the edited video completely changed in visual appearance compared to the Source? 2) Textual alignment: How well does the text description align with the edited video?
Figure 14: Visualization of our stylization process within the image-to-3D pipeline. The source image prompts chosen for this demonstration were the subjects ‘pig’ and ‘crab’. Across these distinct subjects, our method maintained robust editing performance, showcasing its consistency in 3D stylization.